# Augmented Reality GPS Global Positioning System

### v1

@@ Abstract.
Embedded systems development is a complex task that often requires team collaboration. Given the growing market of freelancers and the global shift to remote work, remote collaboration has become a necessity for many developers and clients. While existing communication and coordination tools help users share, discuss, and edit code collaboratively, these tools were specifically designed for software rather than hardware development. In this work, our goal is to explore the design space of remote support tools for embedded systems development. To do this, we interviewed 12 seasoned embedded systems developers regarding their current remote work practices, issues, and needs. We then conducted a user enactment study with a bespoke remote manipulation agent, Handy, as a hypothetical assistant to elicit the types of support developers desire from a collaborator. Our findings describe the scenarios and strategies in which remote work takes place; the support needs and information, coordination, and implementation challenges expressed by developers; and the privacy, control, and trust concerns that developers have when working on their projects with remote physical manipulation tools. This research contributes to the literature by bringing embedded systems development in line with remote, on-demand collaboration and help-seeking in software environments. The empirical basis of this work provides a rich foundation of documented needs, preferences, and desires that can ground future work on remote manipulation agents and enhance collaboration support in the domain of embedded systems development.

### v2

@@ Abstract.
Developing embedded systems is a complex endeavor that frequently requires collaborative teamwork. With the rise of freelance work and the global shift towards remote work, the need for effective remote collaboration has become crucial for many developers and their clients. However, current communication and coordination tools are predominantly tailored for software development rather than hardware-focused tasks. This study investigates the potential for remote support tools specifically designed for embedded systems development. Through interviews with 12 experienced embedded systems developers, we explored their existing remote work practices, challenges, and requirements. We also conducted a user enactment study featuring a custom-designed remote manipulation agent, Handy, as a theoretical assistant, to identify the kinds of support developers would value in a collaborative setting. Our findings highlight the scenarios and strategies employed in remote work, the specific support needs, and the challenges related to information exchange, coordination, and execution. Additionally, we explore concerns around privacy, control, and trust when using remote physical manipulation tools. This research contributes to the field by integrating the development of embedded systems with the remote, on-demand collaboration and assistance typical of software environments, offering a solid empirical foundation for future research on remote manipulation agents in this area.